# Stochastic Control Methods for Optimization

### v1

@@ 2.2 Well-posedness of the stochastic control problems
    The function $G:\mathbb{R}^{d}\to\mathbb{R}$ is nonnegative and there exist constants $L,M>0$ and $\alpha\in(0,1]$ such that
@@
$e^{-\frac{M}{\varepsilon}}\leq u(t,x)\leq 1,\quad\forall\,(t,x)\in[0,1)\times\mathbb{R}^{d}.$ (2.11)
@@
$\partial_{xx}^{2}u(t,x)$ $=\partial_{xx}^{2}\int_{\mathbb{R}^{d}}e^{-\frac{G(z)}{\varepsilon}}\frac{e^{-\frac{|z-x|^{2}}{2(1-t)}}}{(2\pi(1-t))^{d/2}}\,dz$ (2.17)
$=\int_{\mathbb{R}^{d}}e^{-\frac{G(z)}{\varepsilon}}\partial^{2}_{xx}\left(\frac{e^{-\frac{|z-x|^{2}}{2(1-t)}}}{(2\pi(1-t))^{d/2}}\right)dz$
@@
Therefore, the unique strong solution may be extended to the whole interval $[0,1]$.
@@
It is straightforward to check that the function $V_{\varepsilon}$ defined through the (inverse) Cole-Hopf transformation in (1.11) is a classical solution to the HJB equation (1.7). Denote by $\bar{V}_{\varepsilon}(t,x)$ the value function defined in (1.6). Below, we shall verify that $V_{\varepsilon}$ in (1.11) coincides with $\bar{V}_{\varepsilon}$ and that $\theta^{*}_{\varepsilon}$ defined in (1.13) is the optimal control. For each $(t,x)\in[0,1)\times\mathbb{R}^{d}$ and control $\theta\in\Theta$, applying Itô formula gives that for all $\tau\in[t,1)$,
$V_{\varepsilon}(t,x)$ $=\mathbb{E}\left[V_{\varepsilon}(\tau,X_{\tau}^{t,x;\theta})-\int_{t}^{\tau}\left(\partial_{s}V_{\varepsilon}+\frac{1}{2}\Delta V_{\varepsilon}+\theta\cdot\partial_{x}V_{\varepsilon}\right)(s,X_{s}^{t,x;\theta})ds+\int_{t}^{\tau}\partial_{x}V_{\varepsilon}(s,X_{s}^{t,x;\theta})dW_{s}\right]$
@@
Letting $\tau$ tend to $1$, by the dominated convergence theorem, we obtain
@@
$=\mathbb{E}\left[V_{\varepsilon}(1,X_{1}^{t,x;\theta})+\frac{\varepsilon}{2}\int_{t}^{1}|\theta_{s}|^{2}ds\right]$
@@
where the constant $C>0$ depends only on $p,\alpha,L$, and $\mathbb{E}\left[|x_{0}-\xi|^{2}\right]$.
@@ 2.3 Numerical approximations with examples
To validate the proposed method, we consider the following two benchmark functions commonly used in global optimization.
@@
The global minimum is at $x^{*}=(0,\dots,0)$ with $g(x^{*})=0$. In this experiment, we consider a $20$-dimensional instance of the Ackley function and set the number of particles to $N=1000$. The simulation horizon is $T=1.0$, discretized with $M=2001$ time steps. We choose the regularization parameter $\varepsilon=10^{-300}$ and use $S=1000$ Monte Carlo samples at each time step to estimate the drift term. All particles are initialized at $x_{0}=(5,\dots,5)^{\top}$, and the algorithm is executed for $10$ outer iterations. At termination, the empirical mean of the particle system is
$\bar{X}_{T}\approx\,$ $[-0.0122791,\,-0.00456326,\,-0.01121589,\,-0.02288944,\,0.02035992,\,0.03461961,$
$\quad 0.02942655,\,0.00328112,\,-0.01154488,\,-0.0313587,\,-0.00018172,\,-0.00785759,$
$\quad-0.00059935,\,-0.00673907,\,0.0162123,\,-0.01276138,\,-0.00362341,\,-0.03063308,$
$\quad-0.01251183,\,0.00347963].$
The function value at this point is $g(\bar{X}_{T})\approx 0.085852$, which is close to the global minimum value of 0.
@@ 3.1 Finite Particle Approximation
The first source of error comes from approximating the continuous mean-field distribution with a finite number of particles.
First, we recall Lions’ derivative, which is the key tool for analyzing functions defined on the space of probability measures; see [10, 12] for more details. Let $\mathcal{P}_{2}(\mathbb{R}^{d})$ be the space of probability measures on $\mathbb{R}^{d}$ with finite second moments, equipped with the $2$-Wasserstein metric $\mathcal{W}_{2}$. A function $U:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$ is said to be Lions differentiable at $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ if there exists a random variable $\xi$ defined on some probability space $(\Omega,\mathcal{F},\mathbb{P})$ with law $\mathcal{L}(\xi)=\mu$ such that the mapping
@@
The function $\partial_{\mu}U(\mu)$ is called the Lions derivative of $U$ at $\mu$. If the mapping $(\mu,x)\mapsto\partial_{\mu}U(\mu)(x)$ is continuous on $\mathcal{P}_{2}(\mathbb{R}^{d})\times\mathbb{R}^{d}$, then $U$ is said to be continuously Lions differentiable. Higher-order Lions derivatives can be defined similarly; in particular, the second-order Lions derivative $\partial^{2}_{\mu\mu}U(\mu)(x,y)$ is a matrix in $\mathbb{R}^{d\times d}$ for each $(\mu,x,y)\in\mathcal{P}_{2}(\mathbb{R}^{d})\times\mathbb{R}^{d}\times\mathbb{R}^{d}$.
To focus on this aspect, we assume the following smoothness condition on the solution to the Master Equation (1.20).
@@
    it holds that $|\text{Tr}\{\partial^{2}_{\mu\mu}\mathcal{V}_{\varepsilon}(t,\mu,x,y)\}|\leq L_{\varepsilon}<\infty$, for all $(t,\mu,x,y)\in[0,T]\times\mathcal{P}_{2}(\mathbb{R}^{d})\times\mathbb{R}^{d}\times\mathbb{R}^{d}$.
@@
the uniqueness of the classical solution to the Master Equation (1.20) follows from the verification argument. Indeed, applying the Itô-Kunita-Wentzell formula to any classical solution $\tilde{\mathcal{V}}_{\varepsilon}$ of the Master Equation (1.20) yields the fundamental identity: for any control $\theta\in\Theta$,
$J^{\theta}_{\varepsilon}(t,\mu)=\tilde{\mathcal{V}}_{\varepsilon}(t,\mu)+\mathbb{E}\left[\int_{t}^{T}\frac{1}{2\varepsilon}\left|\theta_{s}+\partial_{\mu}\tilde{\mathcal{V}}_{\varepsilon}(s,\mathcal{L}(X^{t,\xi;\theta}_{s}),X^{t,\xi;\theta}_{s})\right|^{2}ds\right].$
@@
and the equality holds when choosing the optimal feedback control $\theta^{*}=-\partial_{\mu}\tilde{\mathcal{V}}_{\varepsilon}(s,\mathcal{L}(X^{t,\xi;\theta^{*}}_{s}),X^{t,\xi;\theta^{*}}_{s})$. Hence, $\tilde{\mathcal{V}}_{\varepsilon}=\mathcal{V}_{\varepsilon}$. ∎
While Assumption 3.1 is stronger than what is required to guarantee the convergence of the normalized particle value function to the mean-field value function (see, e.g., [3, 12, 19, 21, 43] for convergence results under weaker notions of solution), additional regularity is typically indispensable for establishing quantitative rates of convergence; see [10, 12, 16, 23, 31]. Accordingly, in line with the approach of [31], we impose Assumption 3.1 and work under the existence of a classical solution to the master equation to streamline the derivation of an $\mathcal{O}(1/N)$ error estimate. For results on the existence of classical solutions to master equations under suitable structural and smoothness conditions, we refer to [4, 8, 10, 12, 16] for existence results under appropriate structural and smoothness assumptions. To keep the presentation self-contained, we also state below a classical well-posedness result under $C^{2}$-regularity and convexity conditions on $G$. A notable feature of this convex setting is that the resulting bounds on the Lions derivatives can be chosen uniformly in $\varepsilon$.
@@
    (Regularity): $G$ has Lipschitz continuous derivatives $\partial_{\mu}G$, $\partial^{2}_{\mu\mu}G$, and $\partial_{x}\partial_{\mu}G$ and there exists a constant $K_{G}>0$ such that $|\partial^{2}_{\mu\mu}G(\mu)(x,y)|+|\partial_{x}\partial_{\mu}G(\mu)(x)|\leq K_{G}$ for all $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ and $x,y\in\mathbb{R}^{d}$;
@@
    (Convexity): The total Hessian operator $\mathcal{H}_{G}(\mu)$, defined by its action on perturbations $\phi\in L^{2}(\mathbb{R}^{d},\mu;\mathbb{R}^{d})$ via the quadratic form
    $\langle\mathcal{H}_{G}(\mu)\phi,\phi\rangle:=\mathbb{E}\left[\langle\partial_{x}\partial_{\mu}G(\mu)(X)\cdot\phi(X),\phi(X)\rangle\right]+\tilde{\mathbb{E}}\left[\langle\partial^{2}_{\mu\mu}G(\mu)(X,\tilde{X})\cdot\phi(\tilde{X}),\phi(X)\rangle\right],$
    is a positive semi-definite operator on $L^{2}(\Omega)$.
Then the value function $\mathcal{V}_{\varepsilon}$ of the mean-field control problem (1.19) is the unique classical solution to the Master Equation (1.20). Moreover, $\mathcal{V}_{\varepsilon}$ satisfies the following $\varepsilon$-uniform estimates:
$|\partial_{x}\partial_{\mu}\mathcal{V}_{\varepsilon}(t,\mu,x)|\leq 2K_{G},\quad|\partial^{2}_{\mu\mu}\mathcal{V}_{\varepsilon}(t,\mu,x,y)|\leq 2K_{G},$

### v2

@@ 2.2 Well-posedness of the stochastic control problems
    For the function $G:\mathbb{R}^{d}\to\mathbb{R}$, there exist constants $L,M>0$ and $\alpha\in(0,1]$ such that
@@
$e^{-\frac{M}{\varepsilon}}\leq u(t,x)\leq 1,\quad\forall\,(t,x)\in[0,1]\times\mathbb{R}^{d}.$ (2.10)
@@
$\partial_{xx}^{2}u(t,x)$ $=\int_{\mathbb{R}^{d}}e^{-\frac{G(z)}{\varepsilon}}\partial^{2}_{xx}\left(\frac{e^{-\frac{|z-x|^{2}}{2(1-t)}}}{(2\pi(1-t))^{d/2}}\right)dz$
@@
Therefore, the unique strong solution may be extended to the whole interval $[0,1]$ and we are ready to present the well-posedness result.
@@
It is straightforward to check that the function $V_{\varepsilon}(t,x)$ defined through the (inverse) Cole-Hopf transformation in (1.11) is continuous on $0,1]\times\mathbb{R}^{d}$ and is a classical solution to the HJB equation (1.7) over $0,1)\times\mathbb{R}^{d}$. Denote by $\bar{V}_{\varepsilon}(t,x)$ the value function defined in (1.6). Below, we shall verify that $V_{\varepsilon}$ in (1.11) coincides with $\bar{V}_{\varepsilon}$ and that $\theta^{*}_{\varepsilon}$ defined in (1.13) is the optimal control. For each $(t,x)\in[0,1)\times\mathbb{R}^{d}$ and control $\theta\in\Theta$, applying Itô formula gives that for all $\tau\in[t,1)$,
$V_{\varepsilon}(t,x)$ $=\mathbb{E}\left[V_{\varepsilon}(\tau,X_{\tau}^{t,x;\theta})-\int_{t}^{\tau}\left(\partial_{s}V_{\varepsilon}+\frac{1}{2}\Delta V_{\varepsilon}+\theta\cdot\partial_{x}V_{\varepsilon}\right)(s,X_{s}^{t,x;\theta})ds-\int_{t}^{\tau}\partial_{x}V_{\varepsilon}(s,X_{s}^{t,x;\theta})dW_{s}\right]$
@@
Letting $\tau$ tend to $1$, by the dominated convergence theorem and the continuity of $V_{\varepsilon}$, we obtain
@@
$=\mathbb{E}\left[G(X_{1}^{t,x;\theta})+\frac{\varepsilon}{2}\int_{t}^{1}|\theta_{s}|^{2}ds\right]$
@@
where the constant $C>0$ depends only on $\alpha,L$, and $\mathbb{E}\left[|x_{0}-\xi|^{2}\right]$.
@@ 2.3 Numerical approximations with examples
To illustrate the proposed method, we present two standard benchmark problems from global optimization.
@@
The global minimum is at $x^{*}=(0,\dots,0)$ with $g(x^{*})=0$. In this experiment, we consider a $20$-dimensional instance of the Ackley function and set the number of particles to $N=1000$. The simulation horizon is $T=1.0$, discretized with $M=2001$ time steps, and the mean-field coupling weight is $\lambda=0.75$. We choose the regularization parameter $\varepsilon=10^{-300}$ and use $S=1000$ Monte Carlo samples at each time step to estimate the drift term. All particles are initialized at $x_{0}=(5,\dots,5)^{\top}$, and the algorithm is executed for $10$ outer iterations. At termination, the empirical mean of the particle system is
$\bar{X}_{T}\approx\,$ $[0.00677943,\,-0.02596362,\,0.02645172,\,-0.01391329,\,0.00043321,\,-0.00798279,$
$\quad-0.02765826,\,-0.00654247,\,-0.00301501,\,0.00804032,\,0.01024422,\,0.01322995,$
$\quad-0.03054515,\,0.01779264,\,-0.00474688,\,-0.00306365,\,-0.02286298,\,-0.00381099,$
$\quad 0.00962246,\,0.00973099],$
which is close to the global minimizer at the origin.
It is worth noting that as our primary emphasis is on the theoretical development and analysis of the proposed stochastic control method, the numerical experiments above are presented solely for illustrative purposes, and the reported parameter choices are not optimized for accuracy or computational efficiency. In practical implementations, in addition to key parameters such as the regularization level $\varepsilon$, the particle number $N$, the Monte Carlo sample size $S$, the number of outer iterations $L$, the coupling weight $\lambda$, and the time discretization level $M$, one may further tune the time horizon $T$, employ more sophisticated Monte Carlo estimators or variance-reduction techniques, and introduce suitable scalings of the driving Wiener process $W$. A comprehensive numerical investigation, including systematic parameter calibration and detailed comparisons with established global optimization algorithms, is left for future work.
@@ 3.1 Finite Particle Approximation
The first source of error comes from approximating the continuous mean-field control problem with controlled finite number of particles.
First, we recall Lions’ derivative, which is the key tool for analyzing functions defined on the space of probability measures; see [10, 12] for more details. Let $\mathcal{P}_{2}(\mathbb{R}^{d})$ be the space of probability measures on $\mathbb{R}^{d}$ with finite second moments, equipped with the $2$-Wasserstein metric $\mathcal{W}_{2}(\cdot,\cdot)$. A function $U:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$ is said to be (Lions) differentiable at $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ if there exists a random variable $\xi$ defined on some probability space $(\Omega,\mathcal{F},\mathbb{P})$ with law $\mathcal{L}(\xi)=\mu$ such that the mapping
@@
The function $\partial_{\mu}U(\mu)$ is called the (Lions) derivative of $U$ at $\mu$. If the mapping $(\mu,x)\mapsto\partial_{\mu}U(\mu)(x)$ is continuous on $\mathcal{P}_{2}(\mathbb{R}^{d})\times\mathbb{R}^{d}$, then $U$ is said to be continuously (Lions) differentiable. Higher-order (Lions) derivatives can be defined similarly; in particular, the second-order (Lions) derivative $\partial^{2}_{\mu\mu}U(\mu)(x)(y)$ is a matrix in $\mathbb{R}^{d\times d}$ for each $(\mu,x,y)\in\mathcal{P}_{2}(\mathbb{R}^{d})\times\mathbb{R}^{d}\times\mathbb{R}^{d}$.
To focus on the study of particle approximations, we assume the following smoothness condition on the solution to the master equation (1.20).
@@
    it holds that $\left|\int_{\mathbb{R}^{d}}\text{Tr}\left(\partial^{2}_{\mu\mu}\mathcal{V}_{\varepsilon}(t,\mu)(y,y)\right)\mu(dy)\right|\leq L_{\varepsilon}<\infty$, for all $(t,\mu)\in[0,1]\times\mathcal{P}_{2}(\mathbb{R}^{d})$.
@@
The uniqueness of the classical solution to the master equation (1.20) follows from the verification argument. Indeed, applying the Itô formula to any classical solution $\tilde{\mathcal{V}}_{\varepsilon}$ of the master equation (1.20) yields the fundamental identity: for any control $\theta\in\Theta$,
$J^{\theta}_{\varepsilon}(t,\mu)=\tilde{\mathcal{V}}_{\varepsilon}(t,\mu)+\mathbb{E}\left[\int_{t}^{T}\frac{1}{2\varepsilon}\left|\varepsilon\,\theta_{s}+\partial_{\mu}\tilde{\mathcal{V}}_{\varepsilon}(s,\mathcal{L}(X^{t,\xi;\theta}_{s}))(X^{t,\xi;\theta}_{s})\right|^{2}ds\right].$
@@
and the equality holds when choosing the optimal feedback control $\theta^{*}_{s}=-\frac{\partial_{\mu}\tilde{\mathcal{V}}_{\varepsilon}(s,\mathcal{L}(X^{t,\xi;\theta^{*}}_{s}))(X^{t,\xi;\theta^{*}}_{s})}{\varepsilon}$. Hence, $\tilde{\mathcal{V}}_{\varepsilon}=\mathcal{V}_{\varepsilon}$. ∎
While Assumption 3.1 is stronger than what is required to guarantee the convergence of the normalized particle value function to the mean-field value function (see, e.g., [3, 12, 19, 21, 42] for convergence results under weaker notions of solution), additional regularity is typically indispensable for establishing quantitative rates of convergence; see [10, 12, 16, 23, 30]. Accordingly, in line with the approach of [30], we impose Assumption 3.1 and work under the existence of a classical solution to the master equation to streamline the derivation of an $\mathcal{O}(1/N)$ error estimate. For results on the existence of classical solutions to master equations under suitable structural and smoothness conditions, we refer to [4, 8, 10, 12, 16] among many others. For completeness, we also present below a well-posedness result under classical $C^{2}$-type regularity and convexity assumptions on $G$. A key point we want to show is that the corresponding a priori estimates for the Lions derivatives of the value function can be taken uniformly in the regularization parameter $\varepsilon$.
@@
    (Regularity): $G$ has Lipschitz continuous derivatives $\partial_{\mu}G(\mu)(x)$, $\partial^{2}_{\mu\mu}G(\mu)(x)(y)$, $\partial_{x}\partial_{\mu}G(\mu)(x)$, $\partial^{2}_{xx}\partial_{\mu}G(\mu)(x)$, $\partial_{x}\partial^{2}_{\mu\mu}G(\mu)(x)(y)$, and $\partial_{y}\partial_{\mu\mu}^{2}G(\mu)(x)(y)$ and there exists a constant $K_{G}>0$ such that $|\partial^{2}_{\mu\mu}G(\mu)(x)(y)|+|\partial_{x}\partial_{\mu}G(\mu)(x)|\leq K_{G}$ for all $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ and $x,y\in\mathbb{R}^{d}$;
@@
    (Convexity): For each $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, the total Hessian operator $\mathcal{H}_{G}(\mu)$, defined by its action on perturbations $\phi\in L^{2}(\mathbb{R}^{d},\mu;\mathbb{R}^{d})$ via the quadratic form
    $\langle\mathcal{H}_{G}(\mu)\phi,\phi\rangle:=\mathbb{E}\left[\langle\partial_{x}\partial_{\mu}G(\mu)(X)\cdot\phi(X),\phi(X)\rangle\right]+{\mathbb{E}}\left[\langle\partial^{2}_{\mu\mu}G(\mu)(X)(\tilde{X})\cdot\phi(\tilde{X}),\phi(X)\rangle\right],$
    with $X,\tilde{X}$ being independent random variables with law $\mu$, is a positive semi-definite operator.
  3. (c)
    (Coercivity) There exists $\kappa>0$ such that
    $\partial_{x}\partial_{\mu}G(\mu)(x)\geq\kappa I_{d},\quad\text{for all }(x,\mu)\in\mathbb{R}^{d}\times\mathcal{P}_{2}(\mathbb{R}^{d}).$
Then the value function $\mathcal{V}_{\varepsilon}$ of the mean-field control problem (1.19) is the unique classical solution to the master equation (1.20) (as in Assumption 3.1). Moreover, $\mathcal{V}_{\varepsilon}$ satisfies the following $\varepsilon$-uniform estimates:
$|\partial_{x}\partial_{\mu}\mathcal{V}_{\varepsilon}(t,\mu)(x)|\leq\sqrt{d}K_{G},\quad|\partial^{2}_{\mu\mu}\mathcal{V}_{\varepsilon}(t,\mu)(x)(y)|\leq\frac{(K_{G})^{2}}{\kappa}+K_{G},$ (3.1)